OpenStudy (anonymous):

Given:

OpenStudy (amistre64):

should we assume this is a traingle, and that you should employ the law of sines?

OpenStudy (anonymous):

Isnt that a^2= b^2+c^c - 2ab( Cos A) Or something in that order.

OpenStudy (amistre64):

thats the law of cosines :) that would be useful if the setup lends itself to it; but we may need the law of sines to begin with

OpenStudy (amistre64):

sinA/a = sinB/b = sinC/c

OpenStudy (anonymous):

Oh sorry.

OpenStudy (amistre64):



OpenStudy (amistre64):

if we used the law of cosines we would have 1 equation and 2 unknowns: c and C

OpenStudy (anonymous):

so we cross multiply sinA , b and Sin B,a

OpenStudy (amistre64):

\[a^2=b^2+c^2-2bc~cosA\] \[a^2-b^2=c^2-2bc~cosA\] c is stuck in a quadratic that may be difficult at the moment to unwrap. yes, cross multiply to find B, then we know C = 180-(A+B) and can find c

OpenStudy (anonymous):

So \[\frac{ \sin(74) }{ 126}=\frac{ sinB }{ 84 }\]

OpenStudy (anonymous):

sinB*126=80.7

OpenStudy (amistre64):

\[c=\frac{2b~cosA\pm\sqrt{4b^2~cos^2A-4(b^2-a^2)}}{2}\] \[c=\frac{2b~cosA\pm2b\sqrt{cos^2A+a^2}}{2}\] \[c=b~cosA\pm b\sqrt{cos^2A+a^2}\] \[c=b~(cosA\pm \sqrt{cos^2A+a^2})\] :)

OpenStudy (amistre64):

hmm, i almost did that ... but i factored out a b^2 from a^2; so it should prolly be\[c=b~(cosA\pm \sqrt{cos^2A+\frac{a^2}{b^2}})\]

OpenStudy (amistre64):

lets do the sine laws, and see if we get about 156.65 :)

OpenStudy (anonymous):

I get c = 158.3

OpenStudy (amistre64):

\[sin^{-1}(\frac{ 84\sin(74) }{ 126})=B\] \[180-(74+sin^{-1}(\frac{ 84\sin(74) }{ 126}))=105.304^o=C\] \[c = \frac{a~sin(C)}{sin(A)}\] \[c = \frac{126~sin(105.304)}{sin(74)}=126.43\]

OpenStudy (anonymous):

is that what you have for c

OpenStudy (amistre64):

i think so, let me try this simpler .... B = arcsin(b sinA/a) = arcsin(84sin(74)/126) = 39.85 .... C = 180 - 74 - 39.85 = 66.15 c = a sin(C)/sin(A) = 126 sin(66.15)/sin(74) = 119.885

OpenStudy (anonymous):

sorry I lost connection, but ill check that one

OpenStudy (anonymous):

119.885 is correct

OpenStudy (anonymous):

But how do you know when to use the law sines or cosines

OpenStudy (anonymous):

?

OpenStudy (amistre64):

the law of sines is simpler, but can give a false reading in one particular setup

OpenStudy (amistre64):

the law of cosine will always give you a correct answer

OpenStudy (anonymous):

oh, so It really doesnt matter but its best to test both.

OpenStudy (amistre64):



OpenStudy (amistre64):

spose for the law of sines we only know A and c and a, we have 2 solutions for C and b

OpenStudy (anonymous):

okay I think I see what your saying.

OpenStudy (amistre64):

simplicity is best, but can lead to a false reading in a particular case.... so just be wary :)

OpenStudy (anonymous):

okay Ill be sure to test my formulas! :)